# Three-Torsion Subgroups and Wild Conductor Exponents of Plane Quartics

### v3

@@ 5.2. Galois groups and $\text{GSp}_{6}(\mathbb{F}_{3})$
If the degree of $F$ is at least $3^{5}$ and $F$ is square free, then the three-torsion points corresponding to the roots of $F$ generate the entire $3$-torsion subgroup and computing the Galois action on the roots is sufficient to determine the Galois representation. This is since there are more roots of $F$ than non-zero vectors in the largest proper subspace of $\mathbb{F}_{3}^{6}$, and so the torsion points corresponding to these roots must span the whole space. The remark about the Galois action comes from the fact that if the Galois action were not fully determined, then there would be two possible images, which would necessarily map to the same root of $F$, giving a double root.
@@
For each conjugacy class of subgroups, $H\leq\text{GSp}_{6}(\mathbb{F}_{3})$, we compute the order of $H$, the sizes of its orbits on $\mathbb{F}_{3}^{6}\setminus\{0\}$, and the permutation representation on this set. Given a permutation group on $n\leq 728$ symbols, we first restrict to subgroups of the same order, which as remarked above, are stored separately. From this table of subgroups, we compare the orbit sizes: for every orbit of the permutation group, there must exist an orbit of $H$ of the same size if $H$ is the image of the local Galois representation. Finally, for those subgroups satisfying this, we further restrict to those abstractly isomorphic to the permutation group. These isomorphisms are not necessarily unique, for example, by composing with conjugation by a fixed element. This is essentially the only failure, as any two isomorphisms disagree by an automorphism of the group. We compute representatives for the outer automorphism group of the permutation group (the group of automorphisms modulo the subgroup of conjugations), as these will give all the isomorphisms where the action of individual elements may differ. For each isomorphism, we compare the orbit sizes for each conjugacy class of permutations, and its image. As before, if this combination of subgroup and isomorphism realise the permutation group as a matrix group, then for every cycle of the permutation, there exists a cycle of the same size in the permutation action of its image in $H$.
@@ 5.3. Inertia and Ramification Groups
The inertia subgroup, $G_{0}$, is the subgroup of $\mathrm{Gal}(\mathbb{Q}_{\ell}(J[3])/\mathbb{Q}_{\ell})$ consisting of elements, $\sigma$, such that $v_{\ell}(x-\sigma(x))\geq 1$ for all $x\in\mathbb{Q}_{\ell}(J[3])$. Its fixed field is the maximal unramified subfield of $\mathbb{Q}_{\ell}(J[3])$. Similarly, we have the ramification groups in the lower numbering, $G_{i}$, $i\in\mathbb{N}$ is the subgroup consisting of elements $v_{\ell}(x-\sigma(x))\geq i+1$. These groups are $\ell$-groups.
The computer algebra package Magma can compute ramification groups, however due to the work-arounds mentioned in the section on computing splitting fields, this option is not available to us. Instead, we compute these subgroups using a variation of the above definitions. For an element, $\sigma$, of the Galois group to belong to $G_{i}$ it is enough to satisfy $v_{\ell}(\alpha-\sigma(\alpha))\geq i+1$, where $\alpha$ generates the field over $\mathbb{Q}_{\ell}$. We use this combined with the isomorphism identified in the previous section to determine the ramification groups and their action on the 3-torsion.
6. Local Conductor Exponents
The Galois action on the $2$ and $3$ torsion subgroup of $J$ are sufficient to completely the determine the wild conductor exponent at any prime as we now explain.
[…]
6.1. Tame Conductor Exponent
[…]
$n_{p,\text{tame}}=2g-\text{dim}V^{I}$.
This is computable directly from an simple normal crossings regular model of the curve over $\mathbb{Z}_{\ell}$, as we can deduce the following invariants:
The tame part of the exponent is equal to $2g-2a-t$, see [2, Chapter 9] for details. Regular models can be computed in principle by taking any model of the curve and performing repeated blowups until it becomes regular. A number of works have made significant progress towards making regular models accessible, see for instance [7]. However, to date, implementations are slow and inefficient. It is for this reason that we shall not discuss the tame part further.
6.2. Wild Conductor Exponent
[…]
Thus for a given curve explicitly knowing the $\mathbb{Q}_{p}$Galois action on $J[2]$ and $J[3]$ is sufficient to determine all $n_{p,\text{wild}}$. In particular, in the case of plane quartics, one can deduce $n_{l,\text{wild}}$ for all $l\neq 3$, from the $3$-torsion polynomial, as explained in the previous section. In practice we only use the $3$-torsion polynomial to determine $n_{2,\text{wild}}$, and compute all other wild conductor exponent from $J[2]$, as this is particularly efficient to compute in the case of plane quartics.
Remark 5.
6.3. Two-Torsion Subgroups of Jacobians of Plane Quartics
To compute $n_{p,\text{wild}}$ for all odd primes $p$, it suffices to understand the action of $\operatorname{Gal}(\bar{\mathbb{Q}}_{p}/\mathbb{Q}_{p})$ on $J[2]$, for all odd primes $p$ dividing the discriminant. Fortunately, in the case of plane quartics, understanding the $2$-torsion subgroup is straightforward due to its connection with the bitangents to the curve. We give a brief overview here. This is a classical subject and detailed study of this can found in many works, including [10, Chapter 5] and [6]. Let $l$ and $k$ be bitangents to the curve, that is lines in $\mathbb{P}^{2}$ intersecting the curve at $2$ points, each with multiplicity $2$. The function on the curve defined by the quotient of the two lines is of the form
@@ 6.3. Two-Torsion Subgroups of Jacobians of Plane Quartics
for some (not necessarily distinct) points $P_{i},Q_{j}\in C$. It is clear that $D_{l,k}$ defines a $2-$torsion point on the Jacobian. In fact, such points generate the entire $2$-torsion subgroup.
@@
Bitangent lines correspond to odd theta characteristics, that is, square roots of the canonical divisor, whose Riemann-Roch space has odd dimension. As in the construction above, one can clear see that the difference of two such divisors defines a two torsion point on the Jacobian. The fact that such differences generate the entire two-torsion subgroup is a classical result, proved for instance in [10, Chapter 5]. ∎
@@
for some $\alpha_{3},\alpha_{4},\alpha_{5}\in\bar{\mathbb{Q}}$. Equating coefficients in the above expression, we obtain a system of five equations in the $\alpha_{1},\ldots\alpha_{5}$, with $-x+\alpha_{1}y+\alpha_{2}$ defining a bitangent to the curve. We compute a Groebner basis of the ideal defined by these equations (using Magma’s implementation of GroebnerBasis()). Note that if one does not obtain all $28$ bitangents via the above system, the above can be repeated working on a different affine patch, or by assuming a different coefficient to be non-zero, or by composing $f$ with a generic element of $\text{PSL}_{2}(\mathbb{Z})$, to obtain an isomorphic curve. In practice, it is this last option that is most viable as a random element of $\mathrm{PSL}_{2}(\mathbb{Z})$ can be chosen repeatedly until a good choice of model is found.
@@
The sum $D_{k_{1},l_{0}}+...+D_{k_{n},l_{0}}$ is equal to $D_{l,l_{0}}$ if and only if there exists a curve through the intersection points of $C$ with $k_{1}$, …, $k_{n}$, and $l$ of degree $\frac{n+1}{2}$ if $n$ is odd, and through $k_{1}$, …, $k_{n}$, $l$ and $l_{0}$ of degree $\frac{n+2}{2}$ if $n$ is even.
@@
If $n$ is odd, then the equality $D_{k_{1},l_{0}}+...+D_{k_{n},l_{0}}=D_{l,l_{0}}$ can be rearranged to $D_{k_{1},l_{0}}+...+D_{k_{i},l_{0}}=D_{l,l_{0}}+D_{k_{i+1},l_{0}}+...+D_{k_{n},l_{0}}$, where $i=\frac{n+1}{2}$. Adding $\frac{n+1}{2}$ times the intersection points of $l_{0}$ with $C$ to both sides shows this is equivalent to the sum of the intersection points of $k_{1}$, …, $k_{i}$ with $C$ being linear equivalent to the sum of the intersection points of $l$, $k_{i+1},...,k_{n}$ with $C$. The existence of such a curve as in the proposition implies this, let the curve be given by $g=0$, then $\frac{g}{k_{1}...k_{i}}$ gives the desired function. Conversely, if $g=0$ is any curve of degree $\frac{n+1}{2}$ passing through the intersection points of $k_{1}$, …, $k_{i}$ with $C$, then $\frac{g}{k_{1}...k_{i}}$ is a section of the sum of the intersection points. Counting the dimension of such curves, less the dimension of the space of multiples of $f$ (the polynomial defining the curve), gives exactly the expected dimension of the Riemann-Roch space. In particular, if the desired linear equivalence holds, a $g$ can be chosen so that it vanishes at the other intersection points.
The argument for even $n$ is essentially identical for odd $n$, except that an additional factor of the intersection points of $C$ with $l_{0}$ appears. ∎
@@ 6.4.1. The Fermat and Klein Quartic
When the explicit generators of the $3$torsion subgroup are known, computing the wild conductor exponent at $2$ is a simple calculation. For the Fermat quartic, using Magma we compute the extension $\mathbb{Q}_{2}\subseteq K$ over which the $3$-torsion generators are defined, and further compute $G=\operatorname{Gal}\left(K/\mathbb{Q}_{2}\right)$ an the ramification subgroups $\{G_{i}\}_{i\geq-1}$. We found that $G_{n}=1$ for all $n\geq 8$ and $|G_{0}|=|G_{1}|=16$, $|G_{2}|=|G_{3}|=4$ and $|G_{4}|=|G_{5}|=|G_{6}|=|G_{7}|=2$. Explicitly taking Galois invariants, we find:
@@
and thus $n_{2,\text{wild}}=6+2\cdot\frac{6}{4}+4\cdot\frac{2}{8}=10$. Observe that as $J[3]^{G_{0}}=\{0\}$, since the 3-adic Tate module maps onto the 3 torsion, $V^{G_{0}}=\{0\}$ and so the tame part of the conductor is 6. Combining this gives the conductor exponent at 2 as $n_{2}=16$. Since the discriminant of the standard model of the Fermat quartic is $2^{40}$, $2$ is the only prime of bad reduction.
@@ 6.4.2. Generic Examples
For the first 100 entries of Sutherland’s database [27] we computed the wild conductor exponents at all odd primes dividing the discriminant using our implementation. We found that all wild exponents were 00, except for the few exceptions recorded in the table below. The computations took approximately 4.5 hours. We believe that our implementation can be improved greatly if one could reduce the reliance on Magma’s local fields commands. In particular, we found the computation of the splitting field of the bitangents to be very computationally expensive as well as determining a basis, due to a reliance on inexact linear algebra. We note that the initial model of the plane quartic does affect the performance of our algorithm, mostly due to the fact that the precision required to compute the bitangents depends on the coefficients of the equations. As our implementation involves a random choice of model, there are cases for which it is particularly slow.

### v4

@@ 5.2. Galois groups and $\text{GSp}_{6}(\mathbb{F}_{3})$
If the degree of $F$ is at least $3^{5}$ and $F$ is square free, then the three-torsion points corresponding to the roots of $F$ generate the entire $3$-torsion subgroup and computing the Galois action on the roots is sufficient to determine the Galois representation. This is since there are more roots of $F$ than non-zero vectors in the largest proper subspace of $\mathbb{F}_{3}^{6}$, and so the torsion points corresponding to these roots must span the whole space.
@@
For each equivalence class of subgroups under conjugation, $H\leq\text{GSp}_{6}(\mathbb{F}_{3})$, we compute the order of $H$, the sizes of its orbits on $\mathbb{F}_{3}^{6}\setminus\{0\}$, and the permutation representation on this set. Given a permutation group on $n\leq 728$ symbols, we first restrict to subgroups of the same order, which as remarked above, are stored separately. From this table of subgroups, we compare the orbit sizes: for every orbit of the permutation group, there must exist an orbit of $H$ of the same size if $H$ is the image of the local Galois representation. Finally, for those subgroups satisfying this, we further restrict to those abstractly isomorphic to the permutation group. These isomorphisms are not necessarily unique, for example, by composing with conjugation by a fixed element. In fact, any two isomorphisms disagree by an automorphism of the group. We compute representatives for the outer automorphism group of the permutation group (the group of automorphisms modulo the subgroup of conjugations), as these will give all the isomorphisms where the action of individual elements may differ. For each isomorphism, we compare the orbit sizes for each conjugacy class of permutations, and its image. As before, if this combination of subgroup and isomorphism realise the permutation group as a matrix group, then for every cycle of the permutation, there exists a cycle of the same size in the permutation action of its image in $H$.
@@ 5.3. Inertia and Ramification Groups
The inertia subgroup, $G_{0}$, is the subgroup of $G=\mathrm{Gal}(\mathbb{Q}_{\ell}(J[3])/\mathbb{Q}_{\ell})$ consisting of elements, $\sigma$, such that $v_{\ell}(x-\sigma(x))\geq 1$ for all $x$ in the ring of integers of $\mathbb{Q}_{\ell}(J[3])$. Its fixed field is the maximal unramified subfield of $\mathbb{Q}_{\ell}(J[3])$. Similarly, we define the ramification groups in the lower numbering,
$G_{i}=\{\sigma\in G\ |v_{\ell}(x-\sigma(x))\geq i+1\ \text{for all}\ x\in\mathcal{O}_{\mathbb{Q}_{\ell}(J[3])}\}$
for all for all non-negative integers $i$. For any real number $u\geq-1$, set $G_{u}=G_{\lceil u\rceil}$ and define $\phi:[-1,\infty)\rightarrow[-1,\infty)$ by
$\phi(u)=\displaystyle\int_{0}^{u}\frac{1}{[G_{0}:G_{t}]}dt$
if $u\geq 0$ and $\phi(u)=u$ if $1\leq u\leq 0$. The ramification groups in upper numbering are defined as
$G^{i}=G_{\phi(i)}$
for all non-negative integers $i$.
The computer algebra package Magma can compute ramification groups, however due to the work-arounds mentioned in the section on computing splitting fields, this option is not available to us. Instead, we compute these subgroups using a variation of the above definitions. For an element, $\sigma$, of the Galois group to belong to $G_{i}$ it is enough to satisfy $v_{\ell}(\alpha-\sigma(\alpha))\geq i+1$, where $\alpha$ generates ring of integers of the splitting field. We use this combined with the isomorphism identified in the previous section to determine the ramification groups and their action on the 3-torsion.
6. Two-Torsion Subgroup and the Local Galois Representation at 22
For plane quartics, $2$-torsion points have a beautiful geometric description, and they are significantly more manageable to compute. We’ll review this here, and with a view towards computing $n_{p,\text{wild}}$ for all odd primes $p$ in the sections which follows. A detailed overview of bitangents and their connection to $2$-torsion can be found in [10, Chapter 5] and [6], and computational examples can be found in [19].
6.1. Bitangents and $2$-torsion
Let $l$ and $k$ be bitangents to the curve, that is lines in $\mathbb{P}^{2}$ intersecting the curve at $2$ points, each with multiplicity $2$. The function on the curve defined by the quotient of the two lines is of the form
@@ 6.1. Bitangents and $2$-torsion
for some (not necessarily distinct) points $P_{i},Q_{j}\in C$. It is clear that $D_{l,k}:=P_{1}+P_{2}-Q_{1}-Q_{2}$ defines a $2$-torsion point on the Jacobian. In fact, such points generate the entire $2$-torsion subgroup.
@@
Bitangent lines correspond to odd theta characteristics, that is, square roots of the canonical divisor, whose Riemann-Roch space has odd dimension. As in the construction above, one can see that the difference of two such divisors defines a two torsion point on the Jacobian. The fact that such differences generate the entire two-torsion subgroup is a classical result, proved for instance in [10, Chapter 5]. ∎
@@
for some $\alpha_{3},\alpha_{4},\alpha_{5}\in\bar{\mathbb{Q}}$. Equating coefficients in the above expression, we obtain a system of five equations in the $\alpha_{1},\ldots\alpha_{5}$, with $x+\alpha_{1}y+\alpha_{2}$ defining a bitangent to the curve. We compute a Groebner basis of the ideal defined by these equations (using Magma’s implementation of GroebnerBasis()). Note that if one does not obtain all $28$ bitangents via the above system, the above can be repeated working on a different affine patch, or by assuming a different coefficient to be non-zero, or by composing $f$ with a generic element of $\text{PSL}_{2}(\mathbb{Z})$, to obtain an isomorphic curve. In practice, it is this last option that is most viable as a random element of $\mathrm{PSL}_{2}(\mathbb{Z})$ can be chosen repeatedly until a good choice of model is found.
6.2. Local Galois Representation at 22
@@
The sum $D_{k_{1},l_{0}}+...+D_{k_{n},l_{0}}$ is equal to $D_{l,l_{0}}$ if and only if there exists a curve (not containing $C$ as an irreducible component) through the intersection points of $C$ with $k_{1}$, …, $k_{n}$, and $l$ of degree $\frac{n+1}{2}$ if $n$ is odd, and through $k_{1}$, …, $k_{n}$, $l$ and $l_{0}$ of degree $\frac{n+2}{2}$ if $n$ is even.
@@
If $n$ is odd, then the equality $D_{k_{1},l_{0}}+...+D_{k_{n},l_{0}}\sim D_{l,l_{0}}$ can be rearranged to $D_{k_{1},l_{0}}+...+D_{k_{n},l_{0}}+D_{l,l_{0}}\sim 0$. Let $D^{\prime}_{l}$ denote the degree 2 divisor giving the intersection points of a bitangent $l$ with $C$. The desired linear equivalence is equivalent to $D^{\prime}_{k_{1}}+...+D^{\prime}_{k_{n}}+D^{\prime}_{l}\sim(n+1)D^{\prime}_{l_{0}}$. As $l_{0}$ is a bitangent, the right hand side is equivalent to $\frac{n+1}{2}K$. Finally, the linear system of $\frac{n+1}{2}K$ is precisely the intersection of the quartic with degree $\frac{n+1}{2}$ curves (not containing $C$), the result follows.
The argument for even $n$ is essentially identical for odd $n$, except the equality $D^{\prime}_{k_{1}}+...+D^{\prime}_{k_{n}}+D^{\prime}_{l}\sim(n+1)D^{\prime}_{l_{0}}$ is replaced with $D^{\prime}_{k_{1}}+...+D^{\prime}_{k_{n}}+D^{\prime}_{l}+D^{\prime}_{l_{0}}\sim(n+2)D^{\prime}_{l_{0}}\sim\frac{(n+2}{2}K$. ∎
@@
7. Local Conductor Exponents
The Galois action on the $2$ and $3$ torsion subgroup of $J$ are sufficient to completely determine the wild conductor exponent at any prime as we now explain.
$n_{p}=\displaystyle\int_{-1}^{\infty}\text{codim}V^{G_{\mathbb{Q}_{p}}^{u}}\ du$
7.1. Tame Conductor Exponent
$n_{p,\text{tame}}=\displaystyle\int_{-1}^{0}\text{codim}V^{G_{\mathbb{Q}_{p}}^{u}}\ du$.
$n_{p,\text{tame}}=2g-\text{dim}V^{I}$.
This is computable directly from a simple normal crossings regular model of the curve over $\mathbb{Z}_{\ell}$, as we can deduce the following invariants:
The tame part of the exponent is equal to 2g2at2g-2a-t, see [2, Chapter 9] (for the relationship between the abelian and toric ranks and the special fibre) and [17] (for the tame conductor formula, which there is given in terms of the unipotent dimension, gatg-a-t). Regular models can be computed in principle by taking any model of the curve and performing repeated blowups until it becomes regular. A number of works have made significant progress towards making regular models accessible, see for instance [7]. However, to date, implementations are slow and inefficient. It is for this reason that we shall not discuss the tame part further.
7.2. Wild Conductor Exponent
$n_{\text{wild}}=\displaystyle\int_{0}^{\infty}\text{codim}V^{G_{\mathbb{Q}_{p}}^{u}}\ du$.
Thus for a given curve explicitly knowing the $\mathbb{Q}_{p}$ -Galois action on $J[2]$ and $J[3]$ is sufficient to determine all $n_{p,\text{wild}}$. In particular, in the case of plane quartics, one can deduce $n_{l,\text{wild}}$ for all $l\neq 3$, from the $3$-torsion polynomial, as explained in the previous section. In practice we only use the $3$-torsion polynomial to determine $n_{2,\text{wild}}$, and compute all other wild conductor exponent from $J[2]$, as this is particularly efficient to compute in the case of plane quartics.
Remark 7.
@@ 7.3.1. The Fermat and Klein Quartic
When the explicit generators of the $3$-torsion subgroup are known, computing the wild conductor exponent at $2$ is a simple calculation. For the Fermat quartic, using Magma we compute the extension $\mathbb{Q}_{2}\subseteq K$ over which the $3$-torsion generators are defined, and further compute $G=\operatorname{Gal}\left(K/\mathbb{Q}_{2}\right)$ and the ramification subgroups $\{G_{i}\}_{i\geq-1}$. We found that $G_{n}=1$ for all $n\geq 8$ and $|G_{0}|=|G_{1}|=16$, $|G_{2}|=|G_{3}|=4$ and $|G_{4}|=|G_{5}|=|G_{6}|=|G_{7}|=2$. Explicitly taking Galois invariants, we find:
@@
and thus $n_{2,\text{wild}}=6+2\cdot\frac{6}{4}+4\cdot\frac{2}{8}=10$. Observe that as $J[3]^{G_{0}}=\{0\}$, since the 3-adic Tate module maps onto the 3 torsion, $V^{G_{0}}=\{0\}$ and so the tame part of the conductor is 6. Combining this gives the conductor exponent at 2 as $n_{2}=16$. Since the discriminant of the standard model of the Fermat quartic is $^{40}, 22$ is the only prime of bad reduction.
@@ 7.3.2. Generic Examples
For the first 100 entries of Sutherland’s database [29] we computed the wild conductor exponents at all odd primes dividing the discriminant using our implementation. We found that all wild exponents were 0, except for the few exceptions recorded in the table below. The computations took approximately 4.5 hours. We believe that our implementation can be improved greatly if one could reduce the reliance on Magma’s local fields commands. In particular, we found the computation of the splitting field of the bitangents to be very computationally expensive as well as determining a basis, due to a reliance on inexact linear algebra. We note that the initial model of the plane quartic does affect the performance of our algorithm, mostly due to the fact that the precision required to compute the bitangents depends on the coefficients of the equations. As our implementation involves a random choice of model, there are cases for which it is particularly slow.
@@
Declarations
Conflict of Interest
The authors state that there are no conflicts of interest related to this work.
Data availability
Data sharing is not applicable to this article, there are not datasets generated. The examples discussed in this article and the accompanying code can be found in the online repository
https://github.com/ElviraLupoian/Genus3Conductors